OpenStudy (anonymous):

Anyone know how to determine all values of p for which the integral is improper. integral dx/[(x+1)^p] from -2 to 1 ?????????

OpenStudy (anonymous):

it is always improper since the integrand is undefined at \(x=-1\)

OpenStudy (anonymous):

it's discontinuous at x=-1

OpenStudy (anonymous):

yes, so it is always improper question might be for what \(p\) is \[\int_{-2}^1\frac{dx}{(x+1)^p}\] finite

OpenStudy (anonymous):

so p could be all real #s?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

you have to find \[\lim_{t\to -1}\int_{-2}^t\frac{dx}{(x+1)^p}\] to see if you get a finite answer a hint is it will depend on whether \(p>1\) or \(p<1\) try it for \(p=2\) and for \(p=\frac{1}{2}\)

OpenStudy (anonymous):

anti derivative of \(\frac{1}{(x+1)^p}=(x+1)^{-p}\) is \[\frac{(x+1)^{-p+1}}{-p+1}=\frac{1-p}{(x+1)^{p-1}}\]

OpenStudy (anonymous):

wow the last row there, r the 2 sides equal? did u just inverse that? i guess p cannot be 1.

OpenStudy (anonymous):

well i made a mistake

OpenStudy (anonymous):

it is \[\frac{1}{(1-p)(x+1)^{p-1}}\]

OpenStudy (anonymous):

and the question comes down to whether, when you replace \(x\) by -1 you get zero in the denominator (doesn't exist) or a zero in the numerator

OpenStudy (anonymous):

which comes down to whether \(1-p <0\) or \(1-p>0\)